07/29/05 12:00 12M I have posted last years 12 meter receiver tsys measurements.  Look here for the information. News
07/28/05 11:30 ARO All data from both the 12 Meter and the SMT have been archived and removed from the telescope computers. If you need access to your 2004-2005 observing season data, please contact the Operations Manager to obtain a copy of your data. All archive data will be delivered in either DVD-R or CD-R format. News
07/05/05 8:00 ARO The recent Deep Impact Comet Bash was a huge success. Look here for the latest news on all ARO involvement. News
06/30/05 14:30 ARO Portions of our involvement with the Deep Impact comet event will be broadcast live from our Tucson offices Sunday night. All three Tucson affiliates will have camera crews in Steward observatory to provide live feeds for their 10:00 PM news shows. Stefanie Milan will be observing remotely, using both telescopes, from room 154. Channel 12 from Phoenix, the NBC affiliate,  will be interviewing Stefanie at 5:45 & 6:45 AM for their morning show. News
06/30/05 14:00 ARO This coming Sunday night, both telescopes of the ARO observatory will be monitoring the Deep Impact of comet Tempel 1. Look here for more info. News
06/29/05 14:00 SMT The Maser was removed from the SMT and will be shipped back to MIT for repairs and calibration. It will be shipped back and installed in time for next Springs VLBI run. News
06/27/05 16:30 ARO The 12 meter tipper is being shipped to Chile today on a loaner basis to do preliminary site testing for the GMT. It will be back at the 12 meter in time for this Falls observing season. News
06/27/05 12:00 12M After replacing just about everything in the 23mm cryogenic system, the receiver is finally cold. Fixed
06/21/05 12:00 ARO We are announcing a Internal call for observing proposals for the Fall 2005 observing season. The deadline is August 1st. If you have any questions please contact Kiriaki Xilouris, kxiluri{at}as{dot}arizona{dot}edu. Click here for details. News
06/13/05 16:00 SMT The Liebert UPS was repaired with a new input filter. Fixed
06/13/05 12:00 SMT A new up to date status report for the Forbes Filter Bank can be found here. News
06/07/05 04:00 SMT Last Friday, 06/03/05, the Liebert UPS system failed while on a MGIO supplied generator. Here is a report on the weekends activities attempting to bypass this UPS and restore the computer system until at which time the main UPS can be repaired by Titan Power. Bad News
05/25/05 16:00 ARO I have posted a copy of the Forbes Filter Bank design report here. News
05/12/05 10:00 ARO The "Rebel Alliance" has struck Steward Observatory. Look here. News
04/18/05 08:00 SMT The moths have returned to the SMT. Here is a picture of the first one... Bad News
04/11/05 19:00 SMT The JT 1.3mmsis receiver was returned back to the downtown labs for further repairs. Ongoing
04/11/05 10:00 SMT The new Forbes Filter Banks where taken to the SMT today. Here are details of this transportation. News
04/07/05 18:00 12M The tipper has, once again, been taken downtown for repairs. Ongoing
03/18/05 12:00 12M The LAPLACE Astrobiology exchange program with the University of Washington was held Friday, Saturday and Sunday. Here is a link to information on this event. News
03/11/05 17:30 ARO We have finally got a 'Observer Comment Sheet" working and ready for use. You can find it here. News
03/10/05 14:00 12M The 225 GHz tipper has been repaired and returned to service. Fixed
02/17/05 15:30 12M I have posted new versions of the 12 Meter Observers Manual in html format. Also available here in pdf form. News
02/14/05 12:30 SMT This last weekend saw another 36" of snow fall on an already good snow pack on Mt. Graham. Here are a few pictures. News
02/09/05 10:00 SMT The 1mmsis JT receiver was transported to Mt. Graham after repairs in the Steward Lab. News
02/02/05 14:00 12M The 225 GHz tipper has been removed from service and taken downtown for repairs. Ongoing
01/28/05 15:00 ARO ARO has completed and submitted it's report to the Radio, Millimeter and Submillimeter Planning Group. Here is a link to our submission. News
01/20/05 14:00 ARO Several members of the ARO organization attended the the AAS meeting the first week of January. Here is a link to posters and photos. News
01/19/05 17:30 SMT The CPU board in the 'apex' computer system was replaced. This repair will hopefully address the spat of spontaneous reboots we were experiencing with this system. Fixed
01/19/05 12:00 SMT Engineers have discovered a broken IR filter in the optics of the 19 Channel bolometer. This hopefully explains the inability of the bolometer to cool down fully. Repairs are underway. Ongoing
01/18/05 09:00 ARO Proposals are currently being accepted for the Spring 2005 observing season which ends June 30, 2005. The deadline for submission is Feb 7th, 2005. Click here for details. News
01/13/05 14:30 SMT It appears that the 19 Channel Bolometer is having troubles getting cold enough to work. Engineers are going to the site on Tuesday the 18th. to look into it. Ongoing
01/10/05 13:00 SMT The version of Aips on Mt. Graham was too old to patch so I installed the latest version instead. Then I applied the patch described below. News
01/08/05 07:30 12M Thanks to Eric Greisen, we have finally fixed Aips so it can read in our raw sdd files. There was a problem reading Intel byte format files. You may patch your version of Aips 31DEC04 by visiting the Aips site here. I will patch the SMT's installation when I get back from the AAS meeting next week. Fixed
01/03/05 14:30 SMT Looks like the 1mm JT receiver will be out of commission until at least January 31, 2005 Ongoing

 Copyright Arizona Radio Observatory.
For problems or questions regarding this web contact [tfolkers{at}email{dot}arizona{dot}edu].
Last updated: 02/09/10.